OpenStudy (_xoxo_hearts_):

What is the value of u + v + w when u = –10, v = 4, and w = 8?

OpenStudy (dan815):

where you see the letters replace with the numbers u+v+w = -10 + 4 + 8 = ?

OpenStudy (loser66):

I know!!!

OpenStudy (loser66):

dan815 = d+a+n+8+1+5

OpenStudy (_xoxo_hearts_):

I believe it's -22

OpenStudy (dan815):

Nope try again

OpenStudy (dan815):

-10+4+8= -10+12 = ?

OpenStudy (dan815):

-10+12 is the same as +12-10

OpenStudy (_xoxo_hearts_):

2

TheSmartOne (thesmartone):

Use a calculator if it's that tough.

TheSmartOne (thesmartone):

correct :p

OpenStudy (_xoxo_hearts_):

thankies ^.^ @dan815

OpenStudy (_xoxo_hearts_):

could you help me with a few more

OpenStudy (dan815):

:) welcome!